An engine raises 100 kg of water through a height of 60 m in 20 sec. What is the power of the engine?(Take g = 10 m /s²)


1
Expert's answer
2021-04-05T11:08:47-0400
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{100\ kg\cdot10\ \dfrac{m}{s^2}\cdot60\ m}{20\ s}=3000\ W=3\ kW."